OpenStudy (steve816):

When there is a problem that states: Express the function in the form f(g(x)) and for example we take this function F(x) = (2x + x^2)^4 Wouldn't it be technically correct to put it as f(x) = x and g(x) = (2x + x^2)^4

OpenStudy (steve816):

Also, I'm having trouble solving these types of problems, as they are very intuitive and kind of like a puzzle. Any help?

OpenStudy (will.h):

They are pretty easy See the example below and if you still have trouble inform me Suppose we have the following example f(x) = 10x + 2 g(x) = 2x +3 We want to derive the following equation f((gx)) What you need to do is simplify bring the required function in parentheses so (2x + 3) and now see that f(x) has a constant of + 2 so you need to add that and a leading coefficient of 10 so you need to distribute that so we have now 10(2x + 3) +2 = f((gx)) Simplify and that's what we get 20x + 32 = f((gx)) Hope that helps.

OpenStudy (holsteremission):

Setting \(f(x)=x\) and \(g(x)=(2x+x^2)^4\) would be the easiest approach. It's certainly correct that \(F(x)=f(g(x))\) but it's probably not the solution the author had in mind. The next obvious choice would probably be to decompose the composition into its "outer-" and "innermost" components. In this case, you could use \(f(x)=x^4\) and \(g(x)=2x+x^2\).